[18 (-1)*+1(x+1)* k

find radius and interval
Answer #1

(K+1) (-1) 3 (x+1)* 3 k K=1 it converges use Ratio test Ito anth nto an (KD ak altijk k AKF) ==-1){t? (x+y)(k+d) (kt) K aikelkt to at (-1) x H) (HI) K(HD KIA ka ko|i +) +L ak K +) akt) ak s 옥 | +1) </ - X) < ) -20 0x=0 let when +k+1) S (D) kel k = converge at x= -2 o 1-okt!! (-ok Kal k ak+ QO S -K=1 (l k diverges SO - 9xx S 0 ROC = 0-1-2)